Help with HW
the number of tickets sold each day for an upcoming performance of Handel's Messiah is given by N(x)=-0.4x^2+12x +15, where x is the number of days since the concert was first announced. When will daily ticket sales peak and how many tickets will be sold that day?

Answer by stanbon(75887)

the number of tickets sold each day for an upcoming performance of Handel's Messiah is given by N(x)=-0.4x^2+12x +15, where x is the number of days since the concert was first announced. When will daily ticket sales peak and how many tickets will be sold that day?
------------
Max occurs at x = -b/(2a) = -12/(-0.8) = 15
---
f(15) = -0.4(15)^2 + 12(15) + 15 = 186 (max # of tickets sold)
=================
Cheers,
Stan H.